OpenStudy (anonymous):

How would I set this up? f(x) = -3x^2-2x+1

sam (.sam.):

what you want to setup? derivative?

OpenStudy (anonymous):

looks set up to me

OpenStudy (anonymous):

yes, the derivative.

sam (.sam.):

-6x-2

OpenStudy (anonymous):

f'(x)=-(6x+2)

OpenStudy (anonymous):

straight forward application of the "power rule" \[\frac{d}{dx}[x^n]=nx^{n-1}\]

OpenStudy (anonymous):

also \[(cf)'=cf'\] for a constant c, and \[(f+g)'=f'+g'\] meaning you can take the derivative term by term